OpenStudy (anonymous):

find the HCF and LCM of 168101597160 and 23814000 in prime factorisation form

OpenStudy (anonymous):

\[\text{HCF}=2^3*3^5*5^1*7^2 \]\[\text{LCM}=2^4*3^6*5^3*7^8 \]